OpenStudy (anonymous):

Determine the minimum electric potential difference (in volts) that must be applied to accelerate an α-particle so that on interaction with a hydrogen atom, a ground state electron will be excited to n = 6. I need help understanding how to set this up. I am not a chemistry major. I am confused.

OpenStudy (anonymous):

I don't get it either and am doing the same course but I somehow accidentally figured out the correct answer: 6.623

OpenStudy (anonymous):

You can use the following to solve the problem: E = q*V E = -K*(1/n1^2 - 1/n2^2) q*V = -K*(1/n1^2 - 1/n2^2) q = +2e K = 8.987e9 Nm^2/C

OpenStudy (anonymous):

6.60625 volts